OpenStudy (anonymous):

What is the sum of the measures of the interior angles of a regular polygon where a single exterior angle measures 36°? Numerical Answers Expected!

OpenStudy (anonymous):

Assuming the regular polygon has n sides, and thus n interior and n exterior angles, the sum the measures of the exterior angles is always 360 degrees, or 36n = 360 => n = 10. The sum of the measures of the interior angles is (n - 2) 180 = (10 - 2) 180 = 8 x 180 = 1440

OpenStudy (anonymous):

so thats the ans

OpenStudy (anonymous):

A regular hexagon and a regular octagon share a side, as shown in the image below. Regular hexagon and regular octagon share a side forming an exterior angle measuring x degrees What is the measure of x, the angle between a side of the hexagon and a side of the octagon?